# A Survey for Deep Reinforcement Learning Based Network Intrusion Detection

### v1

@@ II-D3 Actor-Critic Network
Figure 5: Classic structure of Actor-Critic network.
@@
Actor-Critic [44] is one of the major DRL algorithms. The Actor learns the policy to take action within a given state directly (Hyper-parameters set is denoted as $\theta$), which is different from DQN (indirectly learning by approaching optimal Q value). The critic is utilized to evaluate the value of the policy by approaching an accurate value function using a neural network (Hyper-parameters set is denoted as $\phi$).
@@
Actor and critic are initialized by $\theta_{0}$ and $\phi_{0}$. Then, based on the initial policy $\pi_{\theta_{0}}(a|s)$, actor starts to explore the environment to collecting interaction experience $(s_{t},a_{t},r_{t},s_{t+1})$. Critic, with initial $\phi_{0}$ parametric, aims to evaluate the value of the policy by calculating temporal difference (TD) in Equation 9 using actor’s experience. The smaller $E(\delta_{t})$ is, the better $\pi_{\theta}$ is. Furthermore, $\phi_{0}$ will be updated using $\delta_{t}$ and gradient decent in Equation 10.
$\delta_{t}=r_{t+1}+\gamma V_{\phi_{0}}(s_{t+1})-V_{\phi_{0}}(s_{t})$ (9)
$\phi_{0}\leftarrow\phi_{0}+\alpha\delta_{t}\nabla_{\phi}V_{\phi_{0}}(s_{t})$ (10)
After updating the Critic, the parametric Actor will be upgraded based on the Critic’s opinion $(\delta_{t})$. Then, $\theta_{0}$ will be updated by Equation 8. What is worth mentioning is that we consider $\delta_{t}$ as a approaching value of the advantage function to represent the relative advantage of an action compared to the average strategy (i.e., current policy $\pi_{\theta}$). Once upgrade parametric Actor, another epoch of collecting experience will go on and repeat the steps above.
$\nabla_{\theta}J(\theta)=E[\nabla_{\theta}\log\pi_{\theta}(a|s_{t})\cdot\delta_{t}]$ (11)
@@
Based on the idea of the actor-critic network, many improved versions have been proposed. Advantage actor critic (A2C) [45] is an improved version of traditional A2C methods by using advantage function $A(s,a)$, which evaluates the extra value of choosing a specific action $a$ compared with an average value of a state, to making a better training efficiency. Based on this idea, the update for actor in A2C is followed by Equation 14 .
$A(s,a)=Q(s,a)-V(s)$ (12)
$Q(s,a)\approx r+\gamma V(s_{t+1},\phi)$ (13)
$\nabla_{\theta}(\theta)=E\left[\nabla\log\pi(a_{t}|s_{t})\cdot A(s_{t},a_{t})\right]$ (14)
$L(\phi)=E\left[(r+\gamma V(s_{t+1},\phi)-V(s_{t},\phi))^{2}\right]$ (15)
@@
The aim of Critic in A2C is making sure to predict the expected return given a specific state $s$ following current policy $\pi$. Thus, trying to minimize the difference between prediction of value function and target return is essential to update the parameters of critic, shown in the Equation 15. Asynchronous Advantage Actor-Critic (A3C) [45] is an algorithm that parallelizes the learning process through multi-threading. Each thread independently explores the environment and calculates gradients, and then asynchronously updates the shared global network parameters [46].
@@ II-F Evaluation
Evaluations in RL are primarily concentrated on assessing the performance of agent in particular task or environment [51]. It relies on specific purpose of tasks, goals of agent and any available feedback information. There are 2 mainly used evaluation metrics, cumulative reward or discounted cumulative reward (return), which means total reward that the agent could gain in one episode or within a fixed period. Success Rate measures the rate at which agent reach specific goals or tasks are successfully completed. It is applicable for tasks with clear success criteria, e.g. navigation. Other evaluation metrics could be set up by understanding the final purpose of task as well, it will depend on the scenario where RL applied. Specifically, in network intrusion detection scenarios, accuracy, precision, recall and $F_{1}$ scores are widely utilized, by calculating True Positive (TP), False Negative (FN), False Positive (FP) and True Negative (TN) from confusion matrix. In some cases, Receiver Operating Characteristics (ROC) curve [52] will be used to measure the model performances.
@@ III Tools used for RL based network intrusion
CSLE is a platform designed for developing and testing reinforcement learning agents in network intrusion detection, offering a realistic cyber range environment. It supports integration with methods like dynamic programming, game theory, and optimization, enhancing research in cyber security. PenGym is a framework for training RL agents in penetration testing, compatible with the Gymnasium API. It allows RL agents to perform actions like network scanning and exploitation in controlled environments. AutoPen is an automated penetration testing framework that uses DRL to identify optimal attack paths in both simulated and real networks. It integrates tools like Nmap and Metasploit to execute attacks, allowing users to study penetration testing techniques for educational purposes. NASimEmu is a framework designed for training DRL agents in offensive penetration-testing scenarios, featuring both a simulator and an emulator for seamless deployment. It uses a random generator to create varied network scenarios and supports simultaneous training across multiple scenarios. CLAP is a RL agent based on PPO that performs penetration testing in simulated network environments using the Network Attack Simulator (NASim). The agent is trained to identify and exploit vulnerabilities to gain access to network resources. Cyberwheel is a RL simulation environment designed for training and evaluating autonomous cyber defence models on simulated networks. Built with modularity, it allows users to customize networks, services, host types, and defensive agents through configurable files. Idsgame is a RL environment designed for simulating attack and defence operations within an abstract network intrusion game. Based on a two-player Markov game model, it features attacker and defender agents competing in a simulated network. The environment provides an interface to a partially observed Markov decision process (POMDP), enabling the training, simulation, and evaluation of attack and defence policies. MAB-Malware is an open-source reinforcement learning framework designed to generate adversarial examples for PE malware by modeling the problem as a multi-armed bandit (MAB). Each action-content pair is treated as an independent slot machine with rewards modeled by a Beta distribution, and Thompson sampling is used to balance exploration and exploitation. YAWNING-TITAN (YT) is a graph-based cyber-security simulation environment built to train intelligent agents for autonomous cyber defence operations. It focuses on simplicity, minimal hardware requirements, and is platform-independent, supporting various algorithms and customizable environment settings.
@@ IV DRL in Network Intrusion detection
DRL has shown remarkable capabilities and results in network intrusion detection past last few years [21, 62, 63, 64, 65]. With combining deep learning and RL, researchers can develop highly efficient, automatic learning detection models. These studies highlight DRL’s powerful ability to process high-dimensional data and complex network environments [65]. It can not only improve detection accuracy and sensitivity, but also optimize its performance through a continuous and dynamic learning process, making network intrusion defence more intelligent and automated [21]. These advantages of DRL herald its widespread application and far-reaching influence in future NID research. In the following subsections, we aim to fully investigate current research and applications for last few years and present insightful views on deep reinforcement learning for network intrusion detection.
@@ IV-A Dataset
Many intrusion datasets have been proposed in recent years. Attack or intrusion types are various in different intrusion datasets. In this section, we are going to present the basic information of current widely utilised network intrusion datasets including KDDCUP 99, NSL-KDD CMU-CERT, UNSW-NB15, CIC-IDS-2017, CSE-CIC-IDS2018, CICDDoS2019, LITNET-2020, AWID, MAWIFlow, CICIoT2023, providing essential understanding including distributions of intrusion types and network traffic features in each datasets and how they were been collected or created. In the end, a comprehensive summery for all datasets are presented for any future reference. The introduction of all mentioned intrusion in the following datasets are presented in Figure 6 for reference.
The KDDCUP 99 dataset [66] was created by processing data from the 1998 DARPA [67] intrusion detection challenge. This was achieved using the Mining Audit Data for Automated Models for Intrusion Detection (MADMAID) framework to extract features from the raw tcpdump data. Detailed statistics of the dataset are provided in Table II. The original 1998 dataset was developed by MIT’s Lincoln Laboratory, involving thousands of UNIX machines and hundreds of users. Network traffic was recorded in tcpdump format  [^2] The tcpdump format is used to capture detailed network traffic information, including timestamps, source and destination IP addresses, port numbers, protocol types, and more. over a ten-week period, with the first seven weeks’ data serving as the training set and the remaining three weeks’ data as the testing set. The KDDCUP 99 DARPA dataset is available in two versions: the full dataset and a 10% sample. It includes 41 features and is categorized into five classes: Normal, DoS, Probe, R2L, and U2R.
@@
CMU-CERT [69] is a synthetic dataset firstly proposed by Computer Emergency and Response Team (CERT) division of Carnegie Mellon University (CMU) as an insider threat dataset. Additionally, this dataset has been continuously updated in recent years. However, one significant drawback is the substantial data imbalance [70]. The insider threat dataset is available in several iterations in recent years, with each new version offering enhancements and improvements. Version 4.2 is notably more frequently utilized, as it includes the highest proportion of intrusions relative to normal data. This version comprises 30,602,325 entries in total, of which 7,623 entries are identified as attacks. Consequently, the percentage of intrusions in this version is approximately 0.025%. The dataset encompasses two years’ worth of Lightweight Directory Access Protocol (LDAP)  [^3] Accessing and managing distributed directory information services. logs, which are instrumental in pinpointing the active users within the company at any given moment. The insider threat dataset, involving 70 employees, are based on three specific scenarios.
@@
Figure 6: An introduction to frequently mentioned intrusion terms. And their frequency across all mentioned datasets.
@@ IV-B Modeling workflow
Figure 7 illustrates the current standard process of implementing a network intrusion detection model using DRL. In the first phase a network dataset is obtained, then, by employing a traffic sampling method, it is possible to extract network features and labels which are needed/utilized in the training phase. In this phase, the policy function $\pi_{\theta}(a|s)$ receives network traffic features data as input and predicts potential intrusion types. Based on these predictions, feedback is provided in the form of correct, incorrect, or uncertain detection outcomes. This feedback is further processed through a reward function, which in turn updates the policy function to enhance its detection accuracy. Designing a practical reward function could accelerate the training speed and the aim of training is to let the policy function get the maximum reward. During the testing phase, the trained policy function is applied to real network traffic, producing intrusion detection results. This entire process iteratively updates the policy function, enabling the system to more effectively identify and classify network intrusions in a dynamic network environment.
@@ IV-C Model performance
DRL-based intrusion models have reached many interesting results, some researchers claim they as a state-of-the-art methods among all intrusion detection models. In this section, we are going to introduce the model performances of DRL-based intrusion detection models in different datasets. Table X shows the performance of DRL models on the NSL-KDD dataset in recent years. Models such as Big-IDS, MAFSIDS, and A-DQN demonstrate excellent performance in terms of accuracy and F1 score, with MAFSIDS particularly standing out, achieving an accuracy of 99.10% and an F1 score of 99.10%. Different models also excel in precision and recall, with the DRL+RBFNN model showing a balanced performance across these metrics. Table XI lists the performances of DRL models on the UNSW-NB15 and CMU-CERT datasets. Overall, these models perform well on the UNSW-NB15 dataset, with the DQN model achieving an accuracy of 91.80% and an F1 score of 92.44%. In comparison, the AE-DQN model also performs notably well on the CMU-CERT dataset, with an accuracy of 88.80% and an F1 score of 89.90%. Table XI shows the performance on the CIC-IDS2017 dataset, models like DRL+RBFNN and A-DQN perform excellently across all metrics, with the DRL+RBFNN model achieving an accuracy of 99.70%, and precision and F1 scores of 99.60% and 99.60%, respectively. Table XI also presents the performance on the CIC-IDS2018 dataset. The ID-RDRL model stands out, particularly in recall and F1 score, achieving 100.00% and 96.30%, respectively. The performance on the CIC-IDS2019 dataset is also shown in Table XI. The ADQN model excels in all metrics, especially in accuracy (99.60%) and F1 score (99.40%). The DQN+CNN model also performs well in terms of precision and recall. For the performance on the AWID dataset, models like AE-SAC and SSDQDN show outstanding performance across all metrics, particularly the AE-SAC model, with an accuracy of 98.98%, and precision and F1 scores of 98.96% and 98.92%, respectively.
IV-D Improved network feature engineering

### v2

@@ II-D3 Actor-Critic Network
Figure 5: Classic structure of an Actor-Critic network.
@@
Actor-Critic [53] is one of the major DRL algorithms. The Actor learns the policy to take action within a given state directly (the hyper-parameters set is denoted as $\theta$), which is different from DQN (indirectly learning by approaching optimal Q value). The critic is utilized to evaluate the value of the policy by approaching an accurate value function using a neural network (Hyper-parameters set is denoted as $\phi$).
@@
Actor and critic are initialized by the $\theta_{0}$ and $\phi_{0}$. Then, based on the initial policy $\pi_{\theta_{0}}(a|s)$, actor starts to explore the environment to collect interaction experience $(s_{t},a_{t},r_{t},s_{t+1})$. Critic, with initial $\phi_{0}$ parameters, aims to evaluate the value of the policy by calculating temporal difference (TD) in Equation 9 using actor’s experience. The smaller $\mathbb{E}(\delta_{t})$ is, the better $\pi_{\theta}$ is. Furthermore, $\phi_{0}$ will be updated using $\delta_{t}$ and gradient descent in Equation 10.
$\delta_{t}=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t})$ (9)
$\phi\leftarrow\phi+\alpha_{\phi}\delta_{t}\nabla_{\phi}V_{\phi}(s_{t})$ (10)
After updating the Critic, the Actor parameters are updated based on the TD error $\delta_{t}$, which serves as an estimate of the advantage function. The policy gradient is computed as shown in Equation 11, and the Actor parameters are updated via gradient ascent: $\theta\leftarrow\theta+\alpha_{\theta}\nabla_{\theta}J(\theta)$, where $\alpha_{\theta}$ is the learning rate for the Actor.
$\nabla_{\theta}J(\theta)=\mathbb{E}_{(s_{t},a_{t})\sim\pi_{\theta}}[\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\cdot\delta_{t}]$ (11)
It is worth noting that the TD error $\delta_{t}$ serves as an unbiased estimate of the advantage function, representing the relative benefit of action $a_{t}$ compared to the baseline value $V(s_{t})$.
@@
Based on the idea of the actor-critic network, many improved versions have been proposed. Advantage Actor-Critic (A2C) [76] is an improved version of the basic Actor-Critic method that incorporates the advantage function $A(s,a)$ to reduce variance in gradient estimates. The advantage function evaluates the relative benefit of choosing a specific action $a$ in state $s$ compared to the expected value under the current policy, thereby improving training stability and efficiency. Based on this idea, the update for actor in A2C is followed by Equation 14 .
A(s,a)=Q(s,a)V(s)A(s,a)=Q(s,a)-V(s) (12)
$Q(s_{t},a_{t})\approx r_{t}+\gamma V(s_{t+1};\phi)$ (13)
$\nabla_{\theta}J(\theta)=\mathbb{E}_{(s_{t},a_{t})\sim\pi_{\theta}}\left[\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\cdot A(s_{t},a_{t})\right]$ (14)
$L(\phi)=\mathbb{E}_{(s_{t},r_{t},s_{t+1})\sim\pi_{\theta}}\left[(r_{t}+\gamma V(s_{t+1};\phi)-V(s_{t};\phi))^{2}\right]$ (15)
@@
The aim of Critic in A2C is making sure to predict the expected return given a specific state $s$ following current policy $\pi$ . Thus, trying to minimize the difference between prediction of value function and target return is essential to update the parameters of critic, shown in the Figure 5. Asynchronous Advantage Actor-Critic (A3C) [76] parallelize the learning process across multiple asynchronous threads (workers). Each worker maintains its own copy of the environment and independently collects experience using the shared global network parameters. After computing local gradients over multiple steps, each worker asynchronously updates the shared global parameters using standard gradient descent, without requiring synchronization among workers. This asynchronous approach improves training efficiency and stability through diverse exploration [113].
@@ II-F Evaluation
Evaluations in RL are primarily concentrated on assessing the performance of agent in particular task or environment [46]. It relies on specific purpose of tasks, goals of agent and any available feedback information. There are 2 mainly used evaluation metrics, cumulative reward or discounted cumulative reward (return), which means total reward that the agent could gain in one episode or within a fixed period. Success rate measures how often the agent reaches its specific goals or successfully completes assigned tasks. It is applicable for tasks with clear success criteria, e.g. navigation. Other evaluation metrics could be set up by understanding the final purpose of task as well, it will depend on the scenario where RL applied. Specifically, in network intrusion detection scenarios, accuracy, precision, recall and $F_{1}$ scores are widely utilized, by calculating True Positive (TP), False Negative (FN), False Positive (FP) and True Negative (TN) from confusion matrix. In some cases, Receiver Operating Characteristics (ROC) curve [36] will be used to measure the model performances. The Area Under the ROC Curve (AUC) shown in Equation 20 quantifies overall discriminative ability by integrating the true positive rate over the false positive rate as the decision threshold $t$ varies, where $TPR(t)=\frac{TP(t)}{TP(t)+FN(t)}$ and $FPR(t)=\frac{FP(t)}{FP(t)+TN(t)}$ denote the true positive rate and false positive rate as functions of the decision threshold $t$, respectively.
@@
PrimAITE PrimAITE is a configurable reinforcement learning simulation environment. PrimAITE
@@ III Tools used for RL based network intrusion
CSLE is a platform designed for developing and testing reinforcement learning agents in network intrusion detection, offering a realistic cyber range environment. It supports integration with methods like dynamic programming, game theory, and optimization, enhancing cyber security research. PenGym is a framework for training RL agents in penetration testing, compatible with the Gymnasium API. It allows RL agents to perform actions like network scanning and exploitation in controlled environments. AutoPen is an automated penetration testing framework that uses DRL to identify optimal attack paths in both simulated and real networks. It integrates tools like Nmap [68] and Metasploit[99] to execute attacks, allowing users to study penetration testing techniques for educational purposes. NASimEmu is a framework designed for training DRL agents in offensive penetration-testing scenarios, featuring both a simulator and an emulator for seamless deployment. It uses a random generator to create varied network scenarios and supports simultaneous training across multiple scenarios. CLAP is a RL agent based on PPO that performs penetration testing in simulated network environments using the Network Attack Simulator (NASim). The agent is trained to identify and exploit vulnerabilities to gain access to network resources. Cyberwheel is a RL simulation environment designed for training and evaluating autonomous cyber defence models on simulated networks. Built with modularity, it allows users to customize networks, services, host types, and defensive agents through configurable files. Idsgame is a RL environment designed for simulating attack and defence operations within an abstract network intrusion game. Based on a two-player Markov game model, it features attacker and defender agents competing in a simulated network. The environment provides an interface to a partially observed Markov decision process (POMDP), enabling the training, simulation, and evaluation of attack and defence policies. MAB-Malware is an open-source reinforcement learning framework designed to generate adversarial examples for PE malware by modeling the problem as a multi-armed bandit (MAB). Each action-content pair is treated as an independent slot machine with rewards modeled by a Beta distribution, and Thompson sampling is used to balance exploration and exploitation. YAWNING-TITAN (YT) is a graph-based cyber-security simulation environment built to train intelligent agents for autonomous cyber defence operations. It focuses on simplicity, minimal hardware requirements, and is platform-independent, supporting various algorithms and customizable environment settings.
In the end, PrimAITE (Primary-level AI Training Environment) is a simulation platform designed for training and evaluating AI agents in cyber defense scenarios. The environment features high configurability through YAML files, realistic network traffic simulation, and support for multiple agents with customizable observation and action spaces. It operates at machine speed to enable rapid training cycles and incorporates reinforcement learning reward functions based on threat mitigation and mission success. PrimAITE ships with example scenarios but is designed as a flexible framework that users can extend and reconfigure to address their specific cyber defense training needs.
@@ IV DRL in Network Intrusion detection
DRL applied to NID has shown remarkable capabilities in the past years [92, 5, 102, 134, 25] as demonstrated by the rapid increase presented in Figure 2 . By combining deep learning and RL, researchers can develop highly efficient, automatic learning detection models. These studies highlight DRL’s powerful ability to process high-dimensional data and complex network environments [25]. It can not only improve detection accuracy and sensitivity, but also optimize its performance through a continuous and dynamic learning process, making network intrusion defence more intelligent and automated [92]. These advantages suggest that DRL will play an increasingly important role in future NID research. In the following subsections, we aim to fully investigate current research and applications for last 5 years and present insightful views on deep reinforcement learning for network intrusion detection.
@@ IV-A Dataset
Many intrusion datasets have been proposed in recent years. Attack or intrusion types are various in different intrusion datasets. In this section, we are going to present the basic information of current widely utilised network intrusion datasets including KDDCUP 99, NSL-KDD CMU-CERT, UNSW-NB15, CIC-IDS-2017, CSE-CIC-IDS2018, CICDDoS2019, LITNET-2020, AWID, MAWIFlow, CICIoT2023, providing essential understanding including distributions of intrusion types and network traffic features in each datasets and how to collect or create them. In the end, a comprehensive summary for all datasets are presented for any future reference. The introduction of all mentioned intrusions in the following datasets is presented in Figure 6 for reference.
The KDDCUP 99 dataset [128] was created by processing data from the 1998 DARPA [74] intrusion detection challenge. This was achieved using the Mining Audit Data for Automated Models for Intrusion Detection (MADMAID) [55, 85] framework to extract features from the raw tcpdump data. Detailed statistics of the dataset are provided in Table II. The original 1998 dataset was developed by MIT’s Lincoln Laboratory, involving thousands of UNIX machines and hundreds of users. Network traffic was recorded in tcpdump format  [^1] The tcpdump format is used to capture detailed network traffic information, including timestamps, source and destination IP addresses, port numbers, protocol types, and more. over a ten-week period, with the first seven weeks’ data serving as the training set and the remaining three weeks’ data as the testing set. The KDDCUP 99 DARPA dataset is available in two versions: the full dataset and a 10% sample. It includes 41 features and is categorized into five classes: Normal, DoS, Probe, R2L, and U2R.
@@
CMU-CERT [130] is a synthetic dataset firstly proposed by Computer Emergency and Response Team (CERT) division of Carnegie Mellon University (CMU) as an insider threat dataset. Additionally, this dataset has been continuously updated in recent years. However, one significant drawback is the substantial data imbalance [117]. The insider threat dataset is available in several iterations in recent years, with each new version offering enhancements and improvements. Version 4.2 is notably more frequently utilized, as it includes the highest proportion of intrusions relative to normal data. This version comprises 30,602,325 entries in total, of which 7,623 entries are identified as attacks. Consequently, the percentage of intrusions in this version is approximately 0.025%. The dataset encompasses two years’ worth of Lightweight Directory Access Protocol (LDAP)  [^2] Accessing and managing distributed directory information services. logs, which are instrumental in pinpointing the active users within the company at any given moment. The insider threat dataset, involving 70 employees, is based on three specific scenarios.
@@
Figure 6: An introduction to frequently mentioned intrusion terms and their frequency across the datasets considered.
@@ IV-B Modeling workflow
Figure 7 illustrates the current standard process of implementing a network intrusion detection model using DRL. In the first phase, a network dataset is obtained, then, by employing a traffic sampling method, it is possible to extract network features and labels which are needed/utilized in the training phase. In this phase, the policy function $\pi_{\theta}(a|s)$ receives network traffic features data as input and predicts potential intrusion types. Based on these predictions, feedback is provided in the form of correct, incorrect, or uncertain detection outcomes. This feedback is further processed through a reward function, which in turn updates the policy function to enhance its detection accuracy. Designing a practical reward function could accelerate the training speed and the aim of training is to let the policy function get the maximum reward. During the testing phase, the trained policy function is applied to real network traffic, producing intrusion detection results. This entire process iteratively updates the policy function, enabling the system to more effectively identify and classify network intrusions in a dynamic network environment.
@@ IV-C Model performance
DRL-based intrusion models have reached many interesting results, some researchers claim they as a state-of-the-art methods among all intrusion detection models. In this section, we are going to introduce the model performances of DRL-based intrusion detection models in different datasets. Table X shows the performance of DRL models on the NSL-KDD dataset in recent years. Models such as Big-IDS, MAFSIDS, and A-DQN demonstrate excellent performance in terms of accuracy and F1 score. Notably, the DQN model proposed by [8] achieves the highest accuracy of 99.36% and F1 score of 99.21% on NSL-KDD dataset, while MAFSIDS also shows strong performance with 99.10% accuracy. It is worth noting that multiple DQN-based models in Table X exhibit significant performance variations (ranging from 81.80% to 99.36%), which can be attributed to differences in network architectures, feature engineering methods, and hyperparameter configurations employed by different researchers. Different models also excel in precision and recall, with the DRL+RBFNN model showing a balanced performance across these metrics. Table XI lists the performances of DRL models on the UNSW-NB15 and CMU-CERT datasets. Overall, these models perform well on the UNSW-NB15 dataset, with the DQN model achieving an accuracy of 91.80% and an F1 score of 92.44%. In comparison, the AE-DQN model also performs notably well on the CMU-CERT dataset, with an accuracy of 88.80% and an F1 score of 89.90%. Table XI shows the performance on the CIC-IDS2017 dataset, models like DRL+RBFNN and A-DQN perform excellently across all metrics, with the DRL+RBFNN model achieving an accuracy of 99.70%, and precision and F1 scores of 99.60% and 99.60%, respectively. Table XI also presents the performance on the CIC-IDS2018 dataset. The ID-RDRL model stands out, particularly in recall and F1 score, achieving 100.00% and 96.30%, respectively. The performance on the CIC-IDS2019 dataset is also shown in Table XI. The ADQN model excels in all metrics, especially in accuracy (99.60%) and F1 score (99.40%). The DQN+CNN model also performs well in terms of precision and recall. For the performance on the AWID dataset, models like AE-SAC and SSDDQN show outstanding performance across all metrics, particularly the AE-SAC model, with an accuracy of 98.98%, and precision and F1 scores of 98.96% and 98.92%, respectively.
IV-D Comparison with Non-DRL Baseline Methods
An important limitation of this survey is that Table X and Table XI exclusively present DRL-based methods without including performance comparisons with non-DRL baseline approaches, which may inadvertently overemphasize the advantages of DRL in network intrusion detection. Recent studies have demonstrated that traditional machine learning algorithms such as XGBoost, Random Forest, and SVM can achieve comparable or even superior performance on benchmark datasets[22, 139]. Specifically, on the NSL-KDD dataset, Random Forest has achieved accuracies ranging from 99.28% to 99.88%, while XGBoost has reached 99.91-99.97% accuracy[121, 86]. And SVM combined with Naïve Bayes has attained 99.35% accuracy[22]. These results are comparable to the best-performing DRL methods reported in Table III (e.g., 99.36% accuracy for DQN). Furthermore, NSL-KDD is increasingly recognized as a relatively simple and outdated dataset that may not adequately represent modern cyber-attacks [22], suggesting that achieving high accuracy on this dataset alone does not necessarily indicate superiority of one approach over another. Future research should incorporate comprehensive comparisons with state-of-the-art non-DRL methods across multiple contemporary datasets to more objectively evaluate the practical advantages of DRL-based approaches in network intrusion detection, particularly in terms of adaptability, real-time learning capabilities, and performance on more complex, realistic datasets that better reflect current threat landscapes.
IV-E Improved network feature engineering